why the electric field due to a spherical shell on its surface is infinity if we calculate using coloumbs law?
 

Grade:12th Pass

1 Answers

suvidhi mehta
6 years ago
i think its not infinity
E = KQ/R^2 using coulombs law [F =KQ1Q2/R^2  OR  E = KQ/R^2]
hope its correct and helps you.